How do I solve this problem (csc x - 1)(csc x + 1)
Answer by chibisan(131)

Since this is not a equation(no equal sign) you can only solve by simplifying(cscx -1)(cscx + 1)
(1/sinx)-1)(1/sinx +1)

(a-b)(a+b) = a^2 - b^2
So,
1/sin^2 x - 1^2
1/sin^2x - 1
1/sin^2x - sin^2x/sin^2x
1-sin^2x/sin^2x

sin^2x+cos^2x = 1
cos^2x = 1-sin^2x

cos^2x/sin^2x
= 1/tan^2x
= cot^2x